OpenStudy (anonymous):

How do I derive cos(4lnt) ??

OpenStudy (anonymous):

My question is differentiate \[t ^{-3}.\cos(4lnt)\] I know it's product rule but I don't have a clue how to 2nd product..

OpenStudy (anonymous):

derive the 2nd product

OpenStudy (amistre64):

its called a chain rule

OpenStudy (amistre64):

essentially you have f(4g(x)) derives to: f'(4(g(x)) * 4g'(x)

OpenStudy (anonymous):

@amistre64 Sorry my internet connection went down. That seems fine. I got (4sin(lnt))/t

OpenStudy (anonymous):

Wolfram says otherwise, can you see where I went wrong.

OpenStudy (anonymous):

I let 4lnt = u

OpenStudy (anonymous):

d/dx cos u = sinu

OpenStudy (amistre64):

t^(-3) * cos(4lnt) t^(-3)' * cos(4lnt) + t^(-3) * cos'(4lnt) -3t^(-4) cos(4lnt) + t^(-3) * (-sin(4lnt))*4ln't -3t^(-4) cos(4lnt) - t^(-3) * sin(4lnt)) * 4/t

OpenStudy (amistre64):

derivatice of cos is -sin

OpenStudy (amistre64):

\[-\frac{cos(4lnt)}{3t^4} - \frac{4sin(4lnt)}{t^3t^1}\] \[-\frac{cos(4lnt)}{3t^4} - \frac{4sin(4lnt)}{t^4}\]

OpenStudy (amistre64):

that 3 should be up top :)

OpenStudy (amistre64):

\[-\frac{1}{t^4}(~3cos(4lnt)+4sin(4lnt)~)\]

OpenStudy (amistre64):

i spose a +C is needed for correctness in the end

OpenStudy (anonymous):

-3t^(-4) cos(4lnt) - t^(-3) * sin(4lnt)) * 4/t Just wondering where the minus on the sin go?

OpenStudy (amistre64):

-3t^(-4) cos(4lnt) - t^(-3) * sin(4lnt)) * 4/t ^^ the -sin part makes the whole thing go negative so it switches the operator

OpenStudy (anonymous):

Oh you brought it out in front of the term.